# On the reconstruction limits of complex networks

### v2

@@ IV.5 Bounds of the information gain
IV.6 Numerical approximations of the mutual information
The mutual information $I(G;X)$ and information gain $\mathcal{I}_{M}(x)$ are generally intractable. Their intractability stems from the evaluation of the posterior, which requires computing of the evidence
$P(X=x)=\sum_{g\in\mathcal{G}}P(G=g)P(X=x|G=g)\,.$ (65)
@@ IV.6 Numerical approximations of the mutual information
IV.6.1 Graph enumeration approach
@@
$\begin{split}I(G;X)\simeq\frac{1}{K}\sum_{k=1}^{K}&\bigg{[}\log P\Big{(}X=x^{(k)}|G=g^{(k)}\Big{)}\\ &-\log P\Big{(}X=x^{(k)}\Big{)}\bigg{]}\,,\end{split}$ (66)
@@
IV.6.2 Variational mean-field approximation
@@
IV.6.3 Graph evidence estimation for the stochastic block model
@@
which is intractable, but can be estimated. In Ref. Peixoto (2021), the author proposes a way to estimate the probability $P(\bm{b}|G)$ of partition given $G$—i.e., the posterior of a Bayesian model for community detection—by sampling a set of $M$ partitions from it using Markov chain Monte Carlo (MCMC). The complete procedure is complex and involves aligning the sampled partitions, identifying aligned partition clusters and estimating the node marginal partition distribution $P(b_{i}=r|G)\equiv\pi_{i,r}(G)$ that node $i$ is in group $r$—we refer to the original paper for technical details. To evaluate the graph marginal log probability, we first notice that
@@
$H_{\mathrm{MF}}(\bm{b}|G)=-\sum_{i=1}^{N}\sum_{r=1}^{B_{\max}}\pi_{i,r}(G)\log\pi_{i,r}(G)$ (74)
where $B_{\max}$ is typically chosen to be equal to $N$, as there can be at most $N$ non-empty groups. Also, note that $H_{\mathrm{MF}}(\bm{b}|G)\geq H(\bm{b}|G)$, since by factoring as in Eq. (73) we assume that the node memberships are conditionally independent, which has the effect of increasing the entropy Cover and Thomas (2006). Finally, the graph evidence entropy can be estimating using that mean-field estimator as follows:
@@
IV.6.4 Evidence estimation for model selection
@@
$\log P(X=x)$ $=\mathbb{E}_{G|X=x}\quantity[\log P(X)]$
$=\mathbb{E}_{G|X=x}\quantity[\log P(X,G)]-H(G|X=x)\,,$ (76)
which follows from the fact that $\log P(X)=\log P(X,G)-\log P(G|X)$. Hence, we build an estimator by sampling from the posterior $M$ graphs $g^{(k)}$ given $x$
$\begin{split}\log P(X=x)\simeq&\frac{1}{M}\sum_{k=1}^{M}\bigg{[}\log P(G=g^{(k)},X=x)\\ &-H(G|X=x)\bigg{]}\,.\end{split}$ (77)
@@
$\begin{split}\log P(X=x)\gtrsim&\frac{1}{M}\sum_{m=1}^{M}\bigg{[}\log P(G=g^{(k)},X=x)\\ &-\sum_{i<i}h\big{(}\pi_{ij}(x)\Big{)}\bigg{]}\,\end{split}$
@@
$\begin{split}\log P(X=x)=\,\,&\mathbb{E}_{\theta,\phi,G|X=x}\quantity[\log P(X,\phi,G,\theta)]\\ &-H(\phi,G,\theta|X=x)\,\end{split}$ (78)
@@
$=H(\phi|X)+H(\theta|G)+H(G|X)\,$ (79)
where $H(\phi|X)=H(\phi|G,\theta,X)$ and $H(\theta|G,X)=H(\theta|G)$, by virtue of the facts that $\phi$ is conditionally independent of $G$ and $\theta$, and that $\theta$ is conditionally independent of $X$. For evaluating $H(\phi|X)$, since $\phi$ are continuous random variables, we estimate the posterior density with kernel density estimation (KDE) with a Gaussian kernel and estimate the differential entropy from the estimated density. In the case of $H(\theta|G)$, this term only concerns the SBM prior in our experiments, where $\theta$ are discrete variables $\bm{b}$ where $b_{i}$ denotes the membership of node $i$ to a group. We use the procedure described in Sec. IV.6.3 to estimate $H(\bm{b}|G)$.
IV.7 Markov chain Monte-Carlo algorithm
@@
This sampling scheme can be generalized when additional hyperparameters $\theta$ of the graph prior or parameters $\phi$ from the likelihood must be inferred as in Sec. II.4. We consider a Gibbs sampling scheme where each random variable $G$, $\theta$ and $\phi$ is sampled sequentially and conditioned on the others. In all cases, the acceptance probability follows Eq. (80), where $G$ is replaced by either $\theta$ or $\phi$ when these parameters are sampled. In this paper, only the SBM among the considered graph models contains parameters of the type of $\theta$. To sample from these, we use the same procedure as in Ref. (Peixoto, 2017, Sec. VI)—we refer to it for further detail. The data models considered contains many parameters that we would like to infer, for example, the infection and recovery probabilities, $\lambda$ and $\beta$, respectively, in the SIS. These parameters are real number constrained within an interval (for instance, $[0,1]$ for the recovery probability $\beta$); Hence, any proposed move where $\varphi$ falls outside of this interval is rejected. We propose moves drawn from a Gaussian distribution with density
@@
In Sec. II.4, we fix $\sigma=0.1$.
IV.8 Reconstructability of graph models with delta distribution
@@
$B=-\sum_{x}q(x)\left[\log p^{*}(x)+\epsilon\left(\frac{q(x)}{p^{*}(x)}-1\right)+\mathcal{O}(\epsilon^{2})\right]$
@@
Also, from the above equation for $H(G)$, we have that the leading terms are $H(G)=\epsilon(\log\epsilon^{-1}+1+\log Z)+\mathcal{O}(\epsilon^{2})$. Consequently, the reconstructability, being the ratio of $I(X;G)$ and $H(G)$, approaches zero as $\epsilon\to 0$ with leading term is $\mathcal{O}\left(\frac{1}{\log\epsilon^{-1}}\right)$.
IV.9 Inference of brain networks
@@
In this section, we describe the procedure we used to reconstruction the mouse brain network from Sec. II.4.1. The raw data is available in Steinmetz et al. (2019), which was originally presented in Ref. Stringer et al. (2019). We refer to their paper for any technical detail regarding the data collection.
@@
Table 3: Statistics for the number of edges determined from the semi-greedy algorithm for each reconstruction model considered in Sec. II.4.1. The highlighted row (SIS with SBM) corresponds to the maximum evidence model associated with Figs. 8 and 9. The average and standard deviations (std. dev.) are obtained from the 8 parallel chains used for the inference.
IV.9.1 Data preprocessing
@@
For the purpose of the experiment, we choose the Krebs recoding which count 1462 monitored neurons. First, we discretize time into $10^{5}$ steps and map each spike time stamp to the correct discrete time interval. Then, since the time duration of the spikes are not available in the original dataset, we artificially extend the spikes for a random duration of $\tau$, which is exponentially distributed with mean $0.012$ seconds—this value corresponds to an approximate activation duration of 10 time steps. Finally, we partition the complete time series into 100 segments of equal size (1000 steps). Figure 6(a) corresponds to the first among the 100 segments of the discretized time series.
IV.9.2 Inference procedure
The inference procedure is very similar to that presented in Appendix IV.7. We consider a model parametrized by $X$, $G$ and their parameters $\phi$ and $\theta$, respectively. However, we have an additional limitation: We do not know the number of edges in the graph. We tried extending our MCMC algorithm by including moves that do not preserve the number of edges—i.e., adding or removing a single edge—, but we found that these attempts suffered from poor mixing time.
@@ IV.9.2 Inference procedure
IV.9.3 Posterior inspection
The graph and parameter marginal posteriors of the maximum evidence model are illustrated in Fig. 8. We also include a validation of the posterior on the inference data. Figure 9 shows the validation, which includes a prediction of the firing rates and the correlation coefficients between connected and disconnected neurons. The inferred graph also allows to reproduce the shape of the cross-correlation density distribution.
@@ Acknowledgments
This work was supported by the Fonds de recherche du Québec – Nature et technologies (SL, VT), the Conseil de recherches en sciences naturelles et en génie du Canada (CM, PD, AA), the Sentinelle Nord initiative funded by the Fonds d’excellence en recherche Apogée Canada (CM, SL, FT, PD, AA), and the Fonds d’accélération des collaboration en santé du Québec – Alliance Neuro-CERVO (PD, AA). We acknowledge Calcul Québec, Digital Research Alliance of Canada and Hectiq.Ai for their technical support and computing infrastructures.

### v3

@@ Appendix E Bounds of the information gain
Appendix F Numerical approximations of the mutual information
The mutual information $I(G;X)$ and information gain $\mathcal{I}_{M}(x)$ are generally intractable. Their intractability stems from the evaluation of the posterior, which requires computing of the evidence, denoted by $\zeta_{M}(x)=P(X=x)$:
$\zeta_{M}(x)=\sum_{g\in\mathcal{G}}P(G=g)P(X=x|G=g)\,.$ (66)
@@ Appendix F Numerical approximations of the mutual information
F.1 Graph enumeration approach
@@
[…]
@@
F.2 Variational mean-field approximation
@@
F.3 Graph evidence estimation for the stochastic block model
@@
which is intractable, but can be estimated. In Ref. Peixoto (2021), the author proposes a way to estimate the probability $P(\bm{b}|G)$ of partition given GG—i.e., the posterior of a Bayesian model for community detection—by sampling a set of $M$ partitions from it using Markov chain Monte Carlo (MCMC). The complete procedure is complex and involves aligning the sampled partitions, identifying aligned partition clusters and estimating the node marginal partition distribution $P(b_{i}=r|G)\equiv\pi_{i,r}(G)$ that node $i$ is in group rr—we refer to the original paper for technical details. To evaluate the graph marginal log probability, we first notice that
@@
$H_{\mathrm{MF}}(\bm{b}|G)=-\sum_{i=1}^{N}\sum_{r=1}^{B_{\max}}\pi_{i,r}(G)\log\pi_{i,r}(G)\,$, (75)
where $B_{\max}$ is typically chosen to be equal to $N$, as there can be at most $N$ non-empty groups. Also, note that $H_{\mathrm{MF}}(\bm{b}|G)\geq H(\bm{b}|G)$, since by factoring as in Eq. (74) we assume that the node memberships are conditionally independent, which has the effect of increasing the entropy Cover and Thomas (2006). Finally, the graph evidence entropy can be estimated using that mean-field estimator as follows:
@@
F.4 Evidence estimation for model selection
@@
$\log\zeta(x)=\mathbb{E}_{G|X=x}\quantity[\log P(X,G)]-H(G|X=x)\,$, (77)
which follows from the fact that $\log P(X)=\log P(X,G)-\log P(G|X)$. Hence, we build an estimator by sampling from the posterior $K$ graphs $g^{(k)}$ given $x$
$\begin{split}\log\zeta(x)\simeq&\frac{1}{K}\sum_{k=1}^{K}\bigg{[}\log P(G=g^{(k)},X=x)\\ &-H(G|X=x)\bigg{]}\,.\end{split}$ (78)
@@
[…]
@@
[…]
@@
$=H(\phi|X)+H(\theta|G)+H(G|X)\,$, (80)
where $H(\phi|X)=H(\phi|G,\theta,X)$ and $H(\theta|G,X)=H(\theta|G)$, by virtue of the facts that $\phi$ is conditionally independent of $G$ and $\theta$, and that $\theta$ is conditionally independent of $X$. For evaluating $H(\phi|X)$, since $\phi$ are continuous random variables, we estimate the posterior density with kernel density estimation (KDE) with a Gaussian kernel and estimate the differential entropy from the estimated density. In the case of $H(\theta|G)$, this term only concerns the SBM prior in our experiments, where $\theta$ are discrete variables $\bm{b}$ where $b_{i}$ denotes the membership of node $i$ to a group. We use the procedure described in Sec. F.3 to estimate $H(\bm{b}|G)$.
Appendix G Markov chain Monte-Carlo algorithm
@@
This sampling scheme can be generalized when additional hyperparameters $\theta$ of the graph prior or parameters $\phi$ from the likelihood must be inferred as in Sec. V. We consider a Gibbs sampling scheme where each random variable $G$, $\theta$ and $\phi$ is sampled sequentially and conditioned on the others. In all cases, the acceptance probability follows Eq. (81), where $G$ is replaced by either $\theta$ or $\phi$ when these parameters are sampled. In this paper, only the SBM among the considered graph models contains parameters of the type of $\theta$. To sample from these, we use the same procedure as in Ref. (Peixoto, 2017, Sec. VI)—we refer to it for further detail. The data models considered contains many parameters that we would like to infer, for example, the infection and recovery probabilities, $\lambda$ and $\beta$, respectively, in the SIS. These parameters are real number constrained within an interval (for instance, $[0,1]$ for the recovery probability $\beta$); Hence, any proposed move where $\varphi$ falls outside of this interval is rejected. We propose moves drawn from a Gaussian distribution with density
@@
In Sec. V, we fix $\sigma=0.1$.
Appendix H Reconstructability of graph models with delta distribution
@@
$B=-\sum_{x}q(x)\left[\log p^{*}(x)+\epsilon\left(\frac{q(x)}{p^{*}(x)}-1\right)+\mathcal{O}(\epsilon^{2})\right]\,.$
@@
Also, from the above equation for $H(G)$, we have that the leading terms are $H(G)=\epsilon(\log\epsilon^{-1}+1+\log Z)+\mathcal{O}(\epsilon^{2})$. Consequently, the reconstructability, being the ratio of $I(X;G)$ and $H(G)$, approaches zero as $\epsilon\to 0$ with leading term $\mathcal{O}\left(\frac{1}{\log\epsilon^{-1}}\right)$.
Appendix I Inference of brain networks
@@
In this section, we describe the procedure we used to reconstruct the mouse brain network from Sec. V.1. The raw data is available in Steinmetz et al. (2019), which was originally presented in Ref. Stringer et al. (2019). We refer to their paper for any technical detail regarding the data collection.
@@
Table 3: Statistics for the number of edges determined from the semi-greedy algorithm for each reconstruction model considered in Sec. V.1. The highlighted row (SIS with SBM) corresponds to the maximum evidence model associated with Figs. 8 and 9. The average and standard deviations (std. dev.) are obtained from the 8 parallel chains used for the inference.
I.1 Data preprocessing
@@
For the purpose of the experiment, we choose the Krebs recoding which count 1462 monitored neurons. First, we discretize time into $10^{5}$ steps and map each spike time stamp to the correct discrete time interval. Then, since the time duration of the spikes are not available in the original dataset, we artificially extend the spikes for a random duration, which is exponentially distributed with mean $0.012$ seconds—this value corresponds to an approximate activation duration of 10 time steps. Finally, we partition the complete time series into 100 segments of equal size (1000 steps). Figure 6(a) corresponds to the first among the 100 segments of the discretized time series.
I.2 Inference procedure
The inference procedure is very similar to that presented in Appendix G. We consider a model parametrized by $X$, $G$ and their parameters $\phi$ and $\theta$, respectively. However, we have an additional limitation: We do not know the number of edges in the graph. We tried extending our MCMC algorithm by including moves that do not preserve the number of edges—i.e., adding or removing a single edge—, but we found that these attempts suffered from poor mixing time.
@@ I.2 Inference procedure
I.3 Posterior inspection
The graph and parameter marginal posteriors of the maximum evidence model are illustrated in Fig. 8. We also include a validation of the posterior on the inference data. Figure 9 shows the posterior predictive check validation, which includes a prediction of the firing rates and the correlation coefficients between connected and disconnected neurons as test quantities. The inferred graph also allows to reproduce the shape of the cross-correlation density distribution.
@@ Acknowledgments
The authors want to thank J.-F. Fortin, N. Doyon and G. Petri for their helpful comments. This work was supported by the Fonds de recherche du Québec – Nature et technologies (SL, VT), the Conseil de recherches en sciences naturelles et en génie du Canada (CM, PD, AA), the Sentinelle Nord initiative funded by the Fonds d’excellence en recherche Apogée Canada (CM, SL, FT, PD, AA), and the Fonds d’accélération des collaboration en santé du Québec – Alliance Neuro-CERVO (PD, AA). We acknowledge Calcul Québec, Digital Research Alliance of Canada and Hectiq.Ai for their technical support and computing infrastructures.